# Simultaneous spatial-parametric collocation approximation for parametric PDEs with log-normal random inputs

### v4

@@ 5 Applications to parametric elliptic PDEs
Let Assumption 5.1 hold. Let the numbers $\alpha$ and $\beta$ be defined by (5.11). Let $\varepsilon$ be an arbitrarily positive number. Then for all $n\geq 2$ there exist points $({\boldsymbol{x}}_{1},{\boldsymbol{y}}_{1}),...,({\boldsymbol{x}}_{n},{\boldsymbol{y}}_{n})\in D\times{\color[rgb]{0,0,0}{{\mathbb{R}}^{\infty}_{0}}}$ and functions $\varphi_{1},...,\varphi_{n}\in V$ and $h_{1},...,h_{n}\in L_{2}({\mathbb{R}}^{\infty},{\mathbb{R}};\gamma)$ such that for the linear sampling algorithm $S_{n}$ on the spatial-parametric domain $D\times{\mathbb{R}}^{\infty}$ defined by
@@
$\begin{split}\big\|u-S_{n}u\big\|_{L_{2}({\mathbb{R}}^{\infty},V;\gamma)}\ \leq\ \begin{cases}Cn^{-\alpha}\quad&{\rm if}\ \alpha\leq 1/p_{2}-3/2,\\ C_{\varepsilon}n^{-\alpha}(\log n)^{\alpha+1+\varepsilon}\quad&{\rm if}\ 1/p_{2}-3/2<\alpha<1/p_{2}-1,\\ C_{\varepsilon}n^{-\alpha}(\log n)^{\alpha+2+\varepsilon}\quad&{\rm if}\ \alpha=1/p_{2}-1,\\ Cn^{-\beta}(\log n)^{\beta+\alpha/(\alpha-\beta)}\quad&{\rm if}\ \alpha>1/p_{2}-1,\end{cases}\end{split}$
for an arbitrarily small number $\varepsilon>0$, where the constants $C$ and $C_{\varepsilon}$ are independent of $n$ .
@@ 6 Extensions of least squares sampling algorithms
As commented in [7], regarding the constructiveness of the linear least squares sampling algorithms in Lemma 6.1, the bound Lemma 6.1(1) is the most coarse bound, but the points construction requires only a random draw, which is computationally inexpensive. The sharper bound in Lemma 6.1(2) uses an additional constructive subsampling step. This was implemented and numerically tested in [8] for up to 1000 basis functions. For larger problem sizes the current algorithm is to slow as its runtime is cubic in the number of basis functions. The sharpest bound in Lemma 6.1(3) is a pure existence result. So, up to now, the only way to obtain this point set is to brute-force every combination, which is computational infeasible.
@@
Let the assumptions of Theorem 2.2 hold. Then for $j=1,2$ and all $n\geq 2$, there exist ${\boldsymbol{y}}_{k,1},...,{\boldsymbol{y}}_{k,n_{k}}\in{\color[rgb]{0,0,0}{{\mathbb{R}}^{\infty}_{0}}}$ and $h_{k,1},...,h_{k,n_{k}}\in L_{2}({\mathbb{R}}^{\infty},{\mathbb{C}};\mu)$, $k=0,...,k_{\lceil n/\log n\rceil}$, such that for the operator $\bar{{\mathcal{S}}}^{X}_{(j),n}$ defined as in (6.1),
@@
$\begin{split}\sup_{v\in B_{X}^{\alpha}}\|v-\bar{{\mathcal{S}}}^{X}_{(1),n}v\|_{L_{2}(U,X;\mu)}\ \leq\ \begin{cases}C_{\varepsilon}n^{-\alpha}(\log n)^{\alpha+1+1/q_{2}+\varepsilon}\quad&{\rm if}\ \alpha<1/q_{2},\\ C_{\varepsilon}n^{-\alpha}(\log n)^{\alpha+2+1/q_{2}+\varepsilon}\quad&{\rm if}\ \alpha=1/q_{2},\\ Cn^{-\beta}(\log n)^{2\beta+\alpha/(\alpha-\beta)}\ \ \quad&{\rm if}\ \alpha>1/q_{2},\end{cases}\end{split}$
for an arbitrarily small number $\varepsilon>0$.
@@
Let the assumptions of Theorem 2.2 hold. Then for $j=1,2$ and all $n\geq 2$, there exist ${\boldsymbol{y}}_{k,1},...,{\boldsymbol{y}}_{k,n_{k}}\in{\color[rgb]{0,0,0}{{\mathbb{R}}^{\infty}_{0}}}$ and $h_{k,1},...,h_{k,n_{k}}\in L_{2}({\mathbb{R}}^{\infty},{\mathbb{C}};\mu)$, $k=0,...,k_{\lceil n/\log n\rceil}$, such that for the operator $\bar{{\mathcal{S}}}^{X}_{(2),n}$ defined as in (6.1),
@@
$\begin{split}\sup_{v\in B_{X}^{\alpha}}\|v-\bar{{\mathcal{S}}}^{X}_{(2),n}v\|_{L_{2}(U,X;\mu)}\ \leq\ C_{\varepsilon}\begin{cases}n^{-\alpha}(\log n)^{\alpha+3/2+\varepsilon}\quad&{\rm if}\ \alpha<1/q_{2},\\ n^{-\alpha}(\log n)^{\alpha+5/2+\varepsilon}\quad&{\rm if}\ \alpha=1/q_{2},\\ n^{-\beta}(\log n)^{3\beta/2+\alpha/(\alpha-\beta)+\varepsilon}\ \quad&{\rm if}\ \alpha>1/q_{2},\end{cases}\end{split}$
for an arbitrarily small number $\varepsilon>0$.
@@
Acknowledgments: This work is funded by the Vietnam National Foundation for Science and Technology Development (NAFOSTED) in the frame of the NAFOSTED–SNSF Joint Research Project under Grant IZVSZ2-229568. A part of this work was done when the author was working at the Vietnam Institute for Advanced Study in Mathematics (VIASM). He would like to thank the VIASM for providing a fruitful research environment and working condition.

### v5

@@ 5 Applications to parametric elliptic PDEs
Let Assumption 5.1 hold. Let $\varepsilon$ be any fixed positive number. Let the numbers $\alpha$ and $\beta$ be defined by (5.9). Then for all $n\geq 2$ there exist points $({\boldsymbol{x}}_{1},{\boldsymbol{y}}_{1}),...,({\boldsymbol{x}}_{n},{\boldsymbol{y}}_{n})\in D\times{\color[rgb]{0,0,0}{{\mathbb{R}}^{\infty}_{0}}}$ and functions $\varphi_{1},...,\varphi_{n}\in V$ and $h_{1},...,h_{n}\in L_{2}({\mathbb{R}}^{\infty},{\mathbb{R}};\gamma)$ such that for the linear sampling algorithm $S_{n}$ on the spatial-parametric domain $D\times{\mathbb{R}}^{\infty}$ defined by
@@
$\big\|u-S_{n}u\big\|_{L_{2}({\mathbb{R}}^{\infty},V;\gamma)}\ \leq\ \begin{cases}Cn^{-\alpha}\quad&{\rm if}\ \alpha\leq{\color[rgb]{0,0,0}{1/p_{2}-1}},\\ C_{\varepsilon}n^{-\alpha}(\log n)^{\alpha+1+\varepsilon}\quad&{\rm if}\ {\color[rgb]{0,0,0}{1/p_{2}-1<\alpha<1/p_{2}-1/2}},\\ C_{\varepsilon}n^{-\alpha}(\log n)^{\alpha+2+\varepsilon}\quad&{\rm if}\ \alpha={\color[rgb]{0,0,0}{1/p_{2}-1/2}},\\ Cn^{-\beta}(\log n)^{\beta+\alpha/(\alpha-\beta)}\quad&{\rm if}\ \alpha>{\color[rgb]{0,0,0}{1/p_{2}-1/2}},\end{cases}$ (5.13)
where the positive constants $C$ depend only on $\alpha,q_{1},q_{2}$ and $K$; the positive constants $C_{\varepsilon}$ depend only on $\alpha,q_{1},q_{2},K,$ and, explicitly, on $\varepsilon$.
Remark 5.5.
We provide commentary on the related results in prior works and contrast them with Theorem 5.4, which presents the main findings of this paper. Recall that the operators $(P_{k})_{k\in{\mathbb{N}}}$ are given by (5.3) and associated with the Lagrange finite element spaces $(V_{k})_{k\in{\mathbb{N}}}$.
We begin by comparing the convergence rates established in Theorem 5.4 with the best-known convergence rates for fully discrete, multilevel Smolyak sparse-grid interpolation. Let $\boldsymbol{I}_{\varepsilon_{n}}^{\operatorname{ML}}$ be the multilevel sparse-grid interpolation algorithm defined in [25, Theorem 7.5] of work at most $n$ which is in particular, based on the approximation properties of $(P_{k})_{k\in{\mathbb{N}}}$ as in Lemma 5.3. Under Assumption 5.1 with the additional condition $p_{2}<3/2$, for the parametric solution $u$ to the equation (1.2) with log-normal random inputs given by (1.3), it has been recently established in [25, Theorem 7.5 and page 177] that for any $n\geq 2$,
$\|u-\boldsymbol{I}_{\varepsilon_{n}}^{\operatorname{ML}}u\|_{L_{2}({\mathbb{R}}^{\infty},V;\gamma)}\ \leq\ C\begin{cases}n^{-\alpha}\log n&\text{if }\alpha\leq 1/p_{2}-3/2,\\ n^{-\beta^{\prime}}\log n&\text{if }\alpha>1/p_{2}-3/2,\end{cases}$ (5.14)
where $\alpha$ is defined in (5.9) and
$\beta^{\prime}:=\left(\frac{1}{p_{1}}-\frac{3}{2}\right)\frac{\alpha}{\alpha+\delta},\quad\delta:=\frac{1}{p_{1}}-\frac{1}{p_{2}}.$
Denote by $a_{n}$ and $b_{n}$ the bounds (without constants) for these convergence rates as in the right-hand sides of (5.13) and (5.14), respectively. Evidently, ignoring values of constants, $a_{n}$ and $b_{n}$ cannot exceed $n^{-\alpha}$ which is governed by the spatial regularity $\alpha$. By simple computation we can see that
$\begin{cases}b_{n}=a_{n}(\log n)^{-1}\ &\text{if }\ \alpha\leq 1/p_{2}-3/2,\\ b_{n}=a_{n}n^{-\tau_{1}}(\log n)^{\alpha+\varepsilon}\ &\text{if }\ 1/p_{2}-3/2<\alpha<1/p_{2}-1/2,\\ b_{n}=a_{n}n^{-\tau_{1}}(\log n)^{\alpha+1+\varepsilon}\ &\text{if }\ \alpha=1/p_{2}-1/2,\\ b_{n}=a_{n}n^{-\tau_{2}}(\log n)^{\beta+\alpha/(\alpha-\beta)-1}\ &\text{if }\ \alpha>1/p_{2}-1/2,\end{cases}$
where
$\tau_{1}:=\alpha-\beta^{\prime}>0\ \ (\text{with}\ \alpha>1/p_{2}-3/2),\qquad\tau_{2}:=\frac{\alpha}{2(\alpha+\delta)}>0.$
This shows that the convergence rate $n^{-\alpha}$ in (5.13) improves the result in (5.14) by a logarithm factor of $(\log n)^{-1}$ in the case of small spatial regularity $\alpha\leq 1/p_{2}-3/2$. The convergence rates in (5.13) significantly improve the convergence rates in (5.14) in the cases of higher spatial regularity $\alpha>1/p_{2}-3/2$. Moreover, in (5.13), the maximal convergence rate $n^{-\alpha}$ still holds with a logarithm factor for the particular intermediate case $1/p_{2}-3/2<\alpha\leq 1/p_{2}-1/2$.
We next compare the established convergence rates in Theorem 5.4 with the best known convergence rates of best $n$-term (and optimal linear) fully discrete approximation applied to the Kondrat’ev analyticity context. Define for $\xi>1$,
$G(\xi):=\ \begin{cases}\big\{(k,{\boldsymbol{s}})\in{\mathbb{N}}_{0}\times{\mathbb{F}}:\,2^{k}\sigma_{2;{\boldsymbol{s}}}^{q_{2}}\leq\xi\big\}\quad&{\rm if}\ \alpha\leq 1/p_{2}-1/2;\\ \big\{(k,{\boldsymbol{s}})\in{\mathbb{N}}_{0}\times{\mathbb{F}}:\,\sigma_{1;{\boldsymbol{s}}}^{q_{1}}\leq\xi,\ 2^{\alpha q_{1}k}\sigma_{2;{\boldsymbol{s}}}^{q_{1}}\leq\xi\big\}\quad&{\rm if}\ \alpha>1/p_{2}-1/2,\end{cases}$
where $q_{j}:=2p_{j}/(2-p_{j})$, $j=1,2$. Denote by ${\mathcal{V}}(G(\xi))$ the subspace in $L_{2}({\mathbb{R}}^{\infty},V;\gamma)$ of all functions $v$ of the form
[…]
We define the linear operator ${\mathcal{S}}_{G(\xi)}:\,L_{2}({\mathbb{R}}^{\infty},{\mathcal{K}}_{\varkappa+1}^{r}(D);\gamma)\to{\mathcal{V}}(G)$ by the truncation of the Hermite GPC expansion
${\mathcal{S}}_{G(\xi)}v:=\ \sum_{(k,{\boldsymbol{s}})\in{G(\xi)}}\delta_{k}(v_{\boldsymbol{s}})H_{\boldsymbol{s}}$
for $v\in L_{2}({\mathbb{R}}^{\infty},{\mathcal{K}}_{\varkappa+1}^{r}(D);\gamma)$ represented by the series (2.39) with $X=V$.
Under Assumption 5.1, we consider the fully discrete (and multi-level) linear approximation by the operator ${\mathcal{S}}_{G(\xi_{n})}$ of the parametric solution $u$ to the equation (1.2) with log-normal random inputs given by (1.3). Building on the results of [20, Theorem 2.1], which establish fully discrete linear approximation in abstract Bochner spaces, together with Lemmata 3.1 and 5.15.4, we obtain the following directly: For each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $\dim({\mathcal{V}}(G(\xi_{n}))\leq n$ and
$\|u-{\mathcal{S}}_{G(\xi_{n})}u\|_{L_{2}({\mathbb{R}}^{\infty},V;\gamma)}\ \leq\ C\begin{cases}n^{-\alpha}&\text{if }\alpha\leq 1/p_{2}-1/2,\\ n^{-\beta}&\text{if }\alpha>1/p_{2}-1/2,\end{cases}$
with the same $\alpha$ and $\beta$ as in Theorem 5.4. As noted in [20, Remark 2.1] these convergence rates agree with the convergence rates of best $n$-term approximation (cf. also [4, Theorem 3.1]). Moreover, the convergence rates stated in Theorem 5.4 coincide with them in the case when $\alpha\leq 1/p_{2}-1$, and differ from them only by a logarithm factor in the other cases.
@@ 6 Extensions of least squares sampling algorithms
As commented in [7], regarding the constructiveness of the linear least squares sampling algorithms in Lemma 6.1, the bound Lemma 6.1(1) is the most coarse bound, but the points construction requires only a random draw, which is computationally inexpensive. The sharper bound in Lemma 6.1(2) uses an additional constructive subsampling step. This was implemented and numerically tested in [8] for up to 1000 basis functions. For larger problem sizes the current algorithm is too slow as its runtime is cubic in the number of basis functions. The sharpest bound in Lemma 6.1(3) is a pure existence result. So, up to now, the only way to obtain this point set is to brute-force every combination, which is computational infeasible.
@@
[…]
@@
$\begin{split}\sup_{v\in B_{X}^{\alpha}}\|v-\bar{{\mathcal{S}}}^{X}_{(1),n}v\|_{L_{2}(U,X;\mu)}\ \leq\ \begin{cases}C_{\varepsilon}n^{-\alpha}{\color[rgb]{0,0,0}{(\log n)^{\alpha+1+\varepsilon}}}\quad&{\rm if}\ \alpha<1/q_{2},\\ C_{\varepsilon}n^{-\alpha}(\log n)^{\alpha+2+1/q_{2}+\varepsilon}\quad&{\rm if}\ \alpha=1/q_{2},\\ Cn^{-\beta}(\log n)^{2\beta+\alpha/(\alpha-\beta)}\ \ \quad&{\rm if}\ \alpha>1/q_{2}.\end{cases}\end{split}$
@@
Let the assumptions of Theorem 2.2 hold. Let $\varepsilon$ be any fixed positive number. Then for $j=1,2$ and all $n\geq 2$, there exist ${\boldsymbol{y}}_{k,1},...,{\boldsymbol{y}}_{k,n_{k}}\in{\color[rgb]{0,0,0}{{\mathbb{R}}^{\infty}_{0}}}$ and $h_{k,1},...,h_{k,n_{k}}\in L_{2}({\mathbb{R}}^{\infty},{\mathbb{C}};\mu)$, $k=0,...,k_{\lceil n/\log n\rceil}$, such that for the operator $\bar{{\mathcal{S}}}^{X}_{(2),n}$ defined as in (6.1),
@@
[…]
@@
Acknowledgments: This work is funded by the Vietnam National Foundation for Science and Technology Development (NAFOSTED) under the Vietnamese–Swiss Joint Research Project, Grant No. IZVSZ2-229568. A part of this work was done when the author was working at the Vietnam Institute for Advanced Study in Mathematics (VIASM). He would like to thank the VIASM for providing a fruitful research environment and working condition.